OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the second equation is substituted into the first equation. 3x + 2y = –21 x – 3y = 4 a. 3(–3y + 4) + 2y = –21 b. 3x + 2(3y + 4) = –21 c. 3(3y + 4) + 2y = –21 d. 3x + 2(–3y + 4) = –21

Parth (parthkohli):

You would first isolate x in the second equation. \( \color{Black}{\Rightarrow x = 4 + 3y}\) Now you will replace \(x\) with \(4 + 3y\).

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3(4 + 3y) + 2y = -21}\)

Parth (parthkohli):

According to the commutative property you can write 4 + 3y as 3y + 4 too \( \color{Black}{\Rightarrow 3(3y + 4) + 2y = -21}\)